OpenStudy (anonymous):

multiply. Give your answer in simplest form. 5 Square root 4x3 square root 21

OpenStudy (anonymous):

\[5\sqrt{4}\times3\sqrt{21}\]

OpenStudy (anonymous):

\[30\sqrt{21}\]

OpenStudy (anonymous):

\[5\sqrt{4}=10 and 10\times3\sqrt{21}=30\sqrt{21}\]

OpenStudy (anonymous):

could you help me with one more

OpenStudy (anonymous):

@elementwielder

OpenStudy (anonymous):

ok whats the question?

OpenStudy (anonymous):

\[\sqrt{\frac{ 33 }{ 144 }}\]

OpenStudy (anonymous):

simplest form?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[\frac{ \sqrt{33} }{ 12 }\]

OpenStudy (anonymous):

33 doesnt have a perfect square root so leave that but square root of 144 is 12

OpenStudy (anonymous):

ok what about \[\frac{ 3\sqrt{20} }{ \sqrt{63} }\]

OpenStudy (anonymous):

\[\frac{ 2\sqrt{35} }{ 7 }\]

OpenStudy (anonymous):

Simplify the radical \[\sqrt{5}+6\sqrt{5}\]

OpenStudy (anonymous):

\[7\sqrt{5}\]

OpenStudy (anonymous):

Simplify the radical \[5\sqrt{7}+2\sqrt{175}\]

OpenStudy (anonymous):

how many more?

OpenStudy (anonymous):

four if your up to it ill fan you if thatd help

OpenStudy (anonymous):

\[15\sqrt{7}\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Simplify the radical \[6\sqrt{5}-8\sqrt{45}\]

OpenStudy (anonymous):

\[-18\sqrt{5}\]

OpenStudy (anonymous):

ok last two

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[\sqrt{26\left( \sqrt{6-3} \right)}\] but the square root only is over 26 not the whole equation

OpenStudy (anonymous):

is that the square root of 6-3?

OpenStudy (anonymous):

its \[\sqrt{6}-3\]

OpenStudy (anonymous):

\[(2\sqrt{3}-3\sqrt{2})\sqrt{13}\]

OpenStudy (anonymous):

\[5=\sqrt{x}-3\]

OpenStudy (anonymous):

x=64

OpenStudy (anonymous):

thanks so much for your help!!

OpenStudy (anonymous):

your welcome

OpenStudy (anonymous):

if you need me to help just ask

OpenStudy (anonymous):

alright will do!

OpenStudy (anonymous):

i kinda need your help some more

OpenStudy (anonymous):

\[\sqrt{5x-1}=\sqrt{4x+9}\]

OpenStudy (anonymous):

@elementwielder

OpenStudy (anonymous):

10

OpenStudy (anonymous):

Solve. \[-5\sqrt{5x-6}=6\sqrt{7x-8}\]

OpenStudy (anonymous):

@elementwielder i could use some help

OpenStudy (anonymous):

\[x=\frac{ 138 }{ 127 }\] if need help visit below website: http://symbolab.com/solver